Q4.

The value of x "\in" R^4

such that (4, 3, - 3, 2) + 3x = (7, -3, 3, 2) is

(1) x = (4, -4, 3, 0)

(2) x = (4, - 4, 3, 0)

(3) x = (1, -2, 0, 0)

(4) None of the given answers is true.


Q5.

Let W be a subset of R^3 defined as

W = (x, y, z) "\in" R^3: 2x + y - z - 1 = 0.

Then

(1) W is a subspace of R^3

(2) W is closed under scalar multiplication

(3) W is not a subspace of R^3

(4) None of the given answers is true.


Q6.

The set of differentiable real-valued functions f on the interval (0,3) such that f'(2) = "\alpha" is a subspace of R^(0,3). The value of must be

(1) negative

(2) positive

(3) zero

(4) None of the given answers is true.


1
Expert's answer
2022-01-25T10:07:04-0500

Q4. "(4,3,-3,2)+3x=(7,-3,3,2)"

"\implies" "3x=(7,-3,3,2)-(4,3,-3,2)"

Subtract component wisely. This gives

"3x=(3,-6,6,0)"

Divide both sides by "3" . This gives

"x=(1,-2,2,0)"

(4)None of the given answers is true.


Q5. Let's check if W is a subspace. First we check if it is closed under addition.

"\begin{bmatrix}
 x_1 \\
 y_1\\
 z_1
\end{bmatrix}" "+\begin{bmatrix}
 x_2 \\
 y_2\\
 z_2
\end{bmatrix}" "=\begin{bmatrix}
 x_1+x_2 \\
 y_1+y_2\\
 z_1+z_2
\end{bmatrix}"

We will test if the points also lies in the plane. So we take our polynomial "2x+y-z-1=0" , and substitute "x" with "x_1+x_2,y" with "y_1+y_2,z" with "z_1+z_2" and get

"2(x_1+x_2)+(y_1+y_2)-(z_1+z_2)-1=0"

When we distribute, we get

"2x_1+2x_2+y_1+y_2-z_1+z_2-1=0"

We reorganize to get

"(2x_1+y_1-z_1)+(2x_2+y_2-z_2)-1=0"

"(2x \n_1\n\u200b\n +y \n_1\n\u200b\n \u2212z \n_1\n\u200b\n )+(2x \n_2\n\u200b\n +y \n_2\n\u200b\n \u2212z \n_2\n\u200b\n )=1"

We cannot say it is closed under addition, since it is not equal to 0.

So

(3) W is not a subspace of "R\u00b3"


Q6. Let "S" be the set of differentiable real valued function. Then

"S=" {"x,f|0<x<3,f\in" R^(0,3) ,"f'(x)=0" }

"\implies f'(x) =0" is a subspace of R^(0,3).

Therefore for "f'(2)=\alpha" to be a subspace, the value of "\alpha" must be zero.

(3) zero